OpenStudy (anonymous):

consider the line y=9x-5 Find the equation of the line that is parallel to this line and passes through the point .(-8,4) Find the equation of the line that is perpendicular to this line and passes through the point (-8,4)

OpenStudy (anonymous):

A line that is parallel will have the same slope, so it will have the form y=9x+c, use the point to find c. Similarly a perpendicular line will have the -1/slope so it will have the form y=-1/9 * x + d

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

ok, lets find c, we input the point in the equation: 4=9(-8)+c, so c=4+9*8= 72, so the first equation is y=9x+72

OpenStudy (anonymous):

do the same for the perpenticular

OpenStudy (anonymous):

so is it y=9x+72?

OpenStudy (anonymous):

the parallel line yes, for the perpendicular you need to find the "d" in y=-1/9 *x +d